OpenStudy (anonymous):

will give medals use the quadratic formula to solve the equation x^2-12=-4x a. 6,2 b. 6,-2 c. -6,2 d. -6,-2

OpenStudy (anonymous):

do you know what the quadratic equation is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x _{1,2} =\frac{ -b \pm \sqrt{b^2-4ac}}{ 2a }\] Does this look familiar to you? Because the question is asking you to use this.

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

do you know how to use it?

OpenStudy (anonymous):

idr

OpenStudy (anonymous):

alright, basically if you have any equation like x^2+7x+9=0 then the number in front of x^2 is a (if there is no number, you use 1) and any number in front of x is b, and the other number is c.

OpenStudy (anonymous):

So in the example I used above it would be a=1, b=7 and c=9. Once you found your a,b and c, you can insert them into the above equation and it should give you two values (one for the + sign and one for the - sign).

OpenStudy (anonymous):

although in your problem you first need to make it into the form ax^2+bx+c=0 because right now it's not like that. But i believe you can do that.

OpenStudy (anonymous):

(1)x^2-4x-12=0?

OpenStudy (anonymous):

If you swap the -4x from one side to the other, remember to change - into + and + into -

OpenStudy (anonymous):

so it's (1)x^2+4x-12?

OpenStudy (anonymous):

yes. See if you add +4x on both sides, the right side will be 0, if you subtract them on both sides then the rght side would be -8x \[(1)x^2\color{green}{+4x}-12=-4x \color{green}{+4x} \]

OpenStudy (anonymous):

ofc -4x + 4x is obviously 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think you can work it out from here, if you need more explanation feel free to ask.

OpenStudy (anonymous):

ok